1. A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?
Show full equation please, I'd like to fully understand it.

Answer by ankor@dixie-net.com(15745)

A plane travels from Orlando to Denver and back again.
On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour.
On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour.
This trip takes six hours. What is the speed of the airplane in still air?
:
Let s = plane speed in still air
then
(s+40) = effective speed with the wind
and
(s-40) = effective speed against
:
The two trips are the same distance, write a distance equation Dist = speed * time
6(s-40) = 5(s+40)
6s - 240 = 5s + 200
6s - 5s = 200 + 240
s = 440 mph in still air
:
:
Check this by finding the actual distances, they should be the same
6*400 = 2400 mi
5*480 = 2400 mi